# Multimodal Laryngoscopic Video Analysis for Assisted Diagnosis of Vocal Fold Paralysis

### v2

@@ Abstract
This paper presents the Multimodal Laryngoscopic Video Analyzing System (MLVAS) [^1] Demo available at https://drive.google.com/file/d/1L1ZH2K3xZiQGFW-1ce89uH9LfYM-RWJH/view?usp=sharing, a novel system that leverages both audio and video data to automatically extract key segments and metrics from raw laryngeal videostroboscopic videos for assisted clinical assessment. The system integrates video-based glottis detection with an audio keyword spotting method to analyze both video and audio data, identifying patient vocalizations and refining video highlights to ensure optimal inspection of vocal fold movements. Additionally, MLVAS features an advanced strobing video extraction module that specifically identifies strobing frames from laryngeal videostroboscopy by analyzing hue, saturation, and value fluctuations. Beyond key segment extraction, MLVAS provides effective metrics for Vocal Fold Paralysis (VFP) detection. It employs a novel two-stage glottis segmentation process using a U-Net for initial segmentation, followed by a diffusion-based refinement to reduce false positives, providing better segmentation masks for downstream tasks. MLVAS estimates the vibration dynamics for both left and right vocal folds from the segmented glottis masks to detect unilateral VFP by measuring the angle deviation with the estimated glottal midline. Comparing the variance between left’s and right’s dynamics, the system effectively distinguishes between left and right VFP. We conducted several ablation studies to demonstrate the effectiveness of each module in the proposed MLVAS. The experimental results on a public segmentation dataset show the effectiveness of our proposed segmentation module. In addition, VFP classification results on a real-world clinic dataset demonstrate MLVAS’s ability of providing reliable and objective metrics as well as visualization for assisted clinical diagnosis.
{IEEEkeywords}
Glottis Segmentation, Keyword Spotting, Laryngoscope, Vocal Fold Paralysis
@@ 1 Introduction
Vocal fold Paralysis (VFP) is a condition where one of the vocal folds fails to move properly, leading to voice changes, difficulty swallowing, and potential breathing problems [1][2]. VFP can result from nerve damage due to surgery, injury, infection, or tumors, significantly impacting a patient’s quality of life [3]. Accurate diagnosis of VFP is crucial, as it informs the appropriate medical or surgical intervention, which can restore vocal function, improve airway protection, and enhance overall patient outcomes. Clinicians often use laryngeal videostroboscopy to check the vocal fold vibration in details. Laryngeal videostroboscopy is a specialized diagnostic tool used to evaluate the function of the vocal folds and the larynx. The stroboscopic component is key because it allows for visualization of vocal fold vibrations, which are often too rapid to be observed directly by doctors [4]. Stroboscopy uses a light source that flashes at a slightly different frequency than the vocal fold vibration. This creates a slow-motion effect, enabling the clinician to see the individual phases of vocal fold vibration in great details.
@@
With the advent of the artificial intelligence, deep learning methods are developed to extract useful parameters [5] to assist clinicians track the motion of vocal folds [6, 7, 8, 9, 10], and even make predictions [11, 12] or classifications [13] to help clinicians perform diagnosis. However, the diagnosis of VFP requires to inspect the complete phonation cycles of patients. Previous works default to using video frames or images that already contain the vocal folds, but in most of the time, the raw recordings from the endoscopic inspection often carry useless information other than patient’s phonation cycles. For instance, no phonation cycles occur in the beginning of the inspection, as the laryngoscope is still finding the vocal folds.
Although experienced experts can provide valuable insights through the analysis of endoscopic videos, this approach often relies on subjective diagnosis, which can compromise objectivity and increase patient anxiety, as well as heighten the risk of misdiagnosis. To address these challenges, various analytical methods have been developed to enhance the efficiency of diagnoses derived from endoscopic videos [14, 15, 16, 17]. Notably, Gloger et al. focus on physical features such as color and shape to segment the glottic area [14], while Adamian et al. and Wang et al. employ supervised learning techniques to track vocal fold movements and derive the Anterior Glottic Angle (AGA), measuring the angle between the two vocal folds [15, 16]. By calculating the glottal area and AGA for each frame, those aforementioned researchers show the Glottal Area Waveform (GAW) and the AGA waveform. However, these methods are limited as those two features essentially measure the gap between left and right vocal folds and therefore cannot indicate which specific vocal fold is paralyzed, rendering them ineffective for diagnosing Unilateral Vocal Fold Paralysis (UVFP). While Pennington et al. expand on feature extraction by including a greater variety of metrics, their methods still derive primarily from the AGA, which limits their applicability for UVFP detection [17]. Additionally, these methods are confined to pre-processed or short video segments without stroboscopic examinations, making the analysis of raw and lengthy laryngoscopic videos in real clinical settings particularly challenging. Consequently, preliminary annotation work is still required in the aforementioned methods, which adds to the inconvenience for clinical experts.
Existing approaches rely solely on video modalities, neglecting the potential benefits of incorporating audio data, which could enhance diagnostic precision. Audio-related methods have shown great potential to detect voice pathology as well [18, 19, 20]. Such methods typically involve converting audio clips to spectrograms using short-time Fourier transform (STFT), which are then fed into various models for downstream tasks. For instance, Deep Neural Networks (DNN) is utilized to predict vocal fold pathology [21, 22, 23, 24]. Low et al. build a machine learning model to detect anomalous patterns in the spectrogram [25]. Compton et al. use an Artificial Neural Network (ANN) to predict vocal fold pathology, claiming to outperform expert assessments [26]. However, these approaches focus solely on the patient’s audio data, without the crucial visual information, which contains more explainable and comprehensive insights into vocal fold conditions.
In this article, we present the Multimodal Laryngoscopic Video Analyzing System (MLVAS), which is a novel system that leverages both the audio and video modalities to automatically process raw laryngeal videostroboscopic videos. The main contributions of our work are listed as follows.
@@
    Multimodal Video Extraction: We propose a system that integrates an audio-based Keyword Spotting (KWS) mechanism to automatically extract valid video segments containing complete phonation cycles and strobing segments from raw laryngeal videostroboscopic videos, effectively reducing the time clinicians spend reviewing lengthy video recordings.
@@
    Two-Stage Segmentation Pipeline: A robust two-stage glottis segmentation framework is developed, utilizing a U-Net model for initial segmentation followed by a diffusion-based refinement stage, which significantly reduces false alarms and enhances segmentation precision.
@@
    Metrics for UVFP Detection: The system introduces Left and Right Vocal Fold Dynamics (LVFDyn and RVFDyn) for UVFP detection. These metrics measure the angle derivation for both sides of the vocal folds through time, which show great potential in UVFP diagnosis.
The organization of this paper is as follows: Section 2 reviews some related works about U-Net based image segmentation and metric computation for UVFP. Section 3 details our proposed method for key video segment extraction, including audio keyword spotting, visual glottis detection, and strobing segment detection. Section 4 and Section 5 focus on the vocal fold analysis module on top of the selected key segments. More specifically, Section 4 presents the two-stage glottis image segmentation technique, including the initial U-Net segmentation and the diffusion based second-pass refinement. Section 5 describes our method to detect VFP and UVFP, including metric computation, glottal midline estimation, and audio-visual VFP detection, respectively. Finally, section 6 presents experimental results, and Section 7 concludes the paper.
2 Related Works
2.1 U-Net Based Glottis Image Segmentation
U-Net’s encoder-decoder structure with skip connections has been proved as a robust framework for biomedical image segmentation [27, 28]. Modifications to this architecture, such as the introduction of separable convolutions in S3AR U-Net [29] and the incorporation of attention mechanisms in Attention U-Net [30], have demonstrated improved performance by capturing more salient features and reducing computational demands. The pursuit of efficiency and clinical applicability has led to the development of models like Efficient U-Net [31], which offers a low inference complexity while maintaining high segmentation quality. Yet, the trade-off between segmentation accuracy and computational cost is a recurring theme, with models like VGG19 U-Net [31] showing longer inference times. Despite the computational efficiency of U-Net variants, the need for substantial labeled training data and the potential for overfitting remain concerns [32]. To counter this, weakly supervised learning approaches, like the one proposed by [33], have gained attraction, requiring only point annotations and demonstrating a remarkable balance between segmentation accuracy and convergence speed.
While the field has made significant strides, challenges persist in achieving a harmonious balance between segmentation accuracy, and generalizability across varied clinical datasets. The limited size of the glottis segmentation mask often leads to a high incidence of false positives when using U-Net structures [34, 35, 36], resulting in segmentation outputs even in the absence of glottal regions. This can significantly impair the accuracy of subsequent analyses. To address this issue, we incorporate a diffusion model as the second-pass refinement, which further enhances the precision of the initial U-Net based segmentation results.
@@ 2.1 U-Net Based Glottis Image Segmentation
2.2 UVFP Detection and glottal midline Estimation
In initial investigations, researchers found that the maximum separation between vocal folds and UVFP are significantly correlated [37]. As studies progressed, the GAW and AGA emerged as key metrics for quantifying this separation, thereby playing a pivotal role in VFP detection. However, existing methodologies typically evaluate those two metrics without distinguishing between the left and right vocal folds [15, 16, 12, 38]. To overcome this limitation, we introduce Vocal Fold Dynamics (VFDyn), including LVFDyn and RVFDyn, assessing the left and right vocal fold angle deviations to the estimated glottal midline. By comparing VFDyn of two vocal folds, our experimental results on actual clinical data indicate that it is possible to effectively differentiate between left and right vocal fold paralysis.
However, assessing the motion of the left and right vocal folds requires accurately identifying the glottal midline, as LVFDyn and RVFDyn represent angle deviations from the midline through time. Traditional methods often depend on extensive labelled datasets for supervised learning [39, 40], which can be labor-intensive. In contrast, our approach autonomously estimates the midline of the glottal area without necessitating any annotated data.
3 Multimodal Key Segment Extraction
Figure 2: The overview of our proposed MLVAS framework.
3.1 System Design
Our system aims to facilitate efficient clinical examinations by extracting key segments from laryngoscopic videos and providing objective indicators for specific laryngeal diseases. As Fig. 2 shows, comprising two main modules – the voice module, the video module – our system ensures the accurate observation of vocalization cycles and the clear visualization of the glottal area.
The voice module initially processes the audio extracted from the video using Short-Time Fourier Transform (STFT) to obtain spectrograms. Through KWS technique, each frame is analyzed to detect patient vocalizations. This enables the preliminary segmentation of vocalization segments within the video.
Subsequently, the video module further refines these vocalization segments to obtain key segment to form laryngoscopic highlights, ensuring the visibility of the vocal folds. Specifically, by utilizing the glottis object detection model, MLVAS can identify regions containing the vocal folds and glottis in each frame. Moreover, given the importance of the stroboscopic portions in laryngoscopic videos for subjective analysis by physicians, we also include a stroboscopic video extraction method into our system.
3.2 Audio Processing Module
To get a better examination of the status of the vocal folds, patients are often pronouncing ”e” (although asked to pronounce ”a”) while the clinicians examine the larynx using the endoscope. In this way, the vibration of the vocal folds can be easily been detected by the clinicians. Hence, by capturing the audio segments that pronounce ”e”, phonation cycles can be detected. To accomplish this, we developed a KWS model that is used to detect audio keywords in a sentence, such as ”Hey Siri”, ”Ok Google”, and in our case, the word ”e”.
The overall pipeline is illustrated in Fig. 1. Initially, the input audio is transformed into a STFT spectrogram, converting the time-domain signal into a time-frequency representation for more effective analysis. The spectrogram is then segmented into chunks along the time axis, with each chunk containing a fixed number of frames. During the training phase, these chunks are randomly selected from each audio clip to ensure a diverse set of training samples. This method enhances the model’s generalization capabilities by incorporating spectrograms from various audio clips into a robust batch of training data.
@@ 3.2 Audio Processing Module
During the inference, the input spectrogram is sliced into chunks using a sliding window mechanism. Each chunk is processed by the trained KWS model to generate decision results, as depicted in Fig. 1. This KWS model can reliably detects the ”ee” vocalization, providing critical prior knowledge for the subsequent analysis.
@@ 3.3 Video Processing Module
Although the audio KWS model is employed to detect the vocalization video segment of the phonation ”e”, these detected video segments may not always successfully capture the vocal folds and glottis on the visual side. To address this issue, we train a vocal fold detection model, utilizing the famous object detection model structure YOLO-v5 [42], to further refine the time masks.
Figure 3: The workflow of bounding box label generation. $U^{\prime},R^{\prime},L^{\prime},D^{\prime}$ are the extended points of $U,R,L,D$ respectively.
Given the limitations of the open-source data for training a detection model, we utilize the public glottis segmentation dataset, BAGLS [43], to construct our training dataset. Since BAGLS is a glottis segmentation dataset, it only contains mask labels for the glottis. To generate labels for vocal fold detection, we designed an automatic bounding box generator. As illustrated in Fig. 3, the process begins by obtaining the coordinates of the top, bottom, left, and right vertices from the glottis mask. These coordinates are denoted as $U(x_{1},y_{1})$, $D(x_{2},y_{2})$, $L(x_{3},y_{3})$, and $R(x_{4},y_{4})$, respectively. Next, these vertices are manually expanded by a fixed number of pixels to ensure adequate coverage around the vocal folds, denoted as $U\prime$, $D\prime$, $L\prime$, and $R\prime$. Using the expanded points, a bounding box is computed, as shown on the right side of Fig. 3.
@@
As depicted in Fig. 2, our pipeline integrates a strobing video extraction module alongside the vocal fold image object detection module. This module is utilized to select the strobing segments of laryngeal videos by analyzing hue, saturation, and value (HSV) of video frames. The HSV parameters represent type, intensity, and brightness of colors, respectively. Transitions marked by empty frames typically occur at the beginning and end of strobing segments. We employ a unit step function to identify and mark all empty frames with a zero value in color (illustrated by the yellow line in Fig. 4). This technique segments the video into several continuous non-empty frame sequences. Within these non-empty segments, we examine the variations in HSV values. The variations is calculated as the total number of changes within a time range. Due to the rapid color-changing characteristic of strobing videos (as shown in Fig. 4), we are able to identify the strobing segments by calculating the frequency of HSV fluctuations within each non-empty segment, selecting the segment with the highest number of fluctuations. Suppose $v_{t-1},v_{t},v_{t+1}$ stands for the value at time step $t-1,t$ and $t+1$ respectively. Then, the frequency of HSV fluctuation within a time range of $T=[t_{0},t_{n}]$ is calculated as follows.
@@
4 Glottis Image Segmentation

### v3

@@ Abstract
This paper presents the Multimodal Laryngoscopic Video Analyzing System (MLVAS) [^1] Demo available at https://drive.google.com/file/d/15on-o7BGB91jwOnz3P2BqzXEhVsTkBjz/view?usp=sharing, a novel system that leverages both audio and video data to automatically extract key video segments and metrics from raw laryngeal videostroboscopic videos for assisted clinical assessment. The system integrates video-based glottis detection with an audio keyword spotting method to analyze both video and audio data, identifying patient vocalizations and refining video highlights to ensure optimal inspection of vocal fold movements. Beyond key video segment extraction from the raw laryngeal videos, MLVAS is able to generate effective audio and visual features for Vocal Fold Paralysis (VFP) detection. Pre-trained audio encoders are utilized to encode the patient voice to get the audio features. Visual features are generated by measuring the angle deviation of both the left and right vocal folds to the estimated glottal midline on the segmented glottis masks. To get better masks, we introduce a diffusion-based refinement that follows traditional U-Net segmentation to reduce false positives. We conducted several ablation studies to demonstrate the effectiveness of each module and modalities in the proposed MLVAS. The experimental results on a public segmentation dataset show the effectiveness of our proposed segmentation module. In addition, unilateral VFP classification results on a real-world clinic dataset demonstrate MLVAS’s ability of providing reliable and objective metrics as well as visualization for assisted clinical diagnosis.
keywords:
Glottis Segmentation, Keyword Spotting, Laryngoscope, Vocal Fold Paralysis, Pre-trained
journal: Computer Speech and Language
[…]
@@ 1 Introduction
Vocal fold Paralysis (VFP) is a condition where one of the vocal folds fails to move properly, leading to voice changes, difficulty swallowing, and potential breathing problems [1, 2]. VFP can result from nerve damage due to surgery, injury, infection, or tumors, significantly impacting a patient’s quality of life [3]. Accurate diagnosis of VFP is crucial, as it informs the appropriate medical or surgical intervention, which can restore vocal function, improve airway protection, and enhance overall patient outcomes. Clinicians often use laryngeal videostroboscopy to check the vocal fold vibration in details. Laryngeal videostroboscopy is a specialized diagnostic tool used to evaluate the function of the vocal folds and the larynx. The stroboscopic component is important because it allows for visualization of vocal fold vibrations, which are often too rapid to be observed directly by doctors [4]. Stroboscopy uses a light source that flashes at a slightly different frequency than the vocal fold vibration. This creates a slow-motion effect, enabling the clinician to see the individual phases of vocal fold vibration in great details.
@@
With the advent of the artificial intelligence, deep learning methods are developed to extract useful parameters [5] to assist clinicians track the motion of vocal folds [6, 7, 8, 9, 10], and even make predictions [11, 12] or classifications [13] to help clinicians perform diagnosis. However, the diagnosis of vocal folds requires to inspect the complete phonation cycles of patients. Previous works default to using video frames or images that already contain the vocal folds, but in most of the time, the raw recordings from the endoscopic inspection often carry useless information other than patient’s phonation cycles. For instance, no phonation cycles occur in the beginning of the inspection, as the laryngoscope is still finding the vocal folds. As a result, manual labeling and selection of video frames from raw endoscopic videos are required, making the diagnosis process time-consuming and inconvenient for clinicians.
Although experienced experts can provide valuable insights through the analysis of endoscopic videos, this approach often relies on subjective diagnosis, which can compromise objectivity and increase patient anxiety, as well as heighten the risk of misdiagnosis. To address these challenges, various analytical methods have been developed to enhance the efficiency of diagnoses derived from endoscopic videos. However, most of the existing approaches rely solely on single modality, audio or video modality, neglecting the potential benefits of combining those two modalities, which could enhance diagnostic precision.
Audio-based methodologies have exhibited remarkable potential in detecting voice pathologies. These techniques often involve transforming audio signals into sophisticated features, such as Mel spectrograms, before further modeling. For instance, Compton et al. proposed a lightweight MLP model for predicting vocal fold pathologies, claiming its performance surpassed expert evaluations [14]. Similarly, Deep Neural Networks (DNNs) have been widely adopted in voice pathology prediction tasks [15, 16, 17, 18]. However, clinical data is often challenging to collect, and datasets are typically restricted from public disclosure due to patient privacy concerns, leading to a scarcity of training samples. Deep learning models, which are prone to overfitting on limited datasets, have driven researchers to use explainable handcrafted features such as Mel-frequency cepstral coefficients (MFCCs) as input for traditional machine learning methods, including SVM, RF, and NB [19, 20, 18].
Another prominent solution for addressing audio data scarcity in specific tasks is the use of pre-trained audio encoders. By fine-tuning these models on small, task-specific datasets or directly use the output feature embeddings of these models, researchers have achieved comparable or even superior performance relative to state-of-the-art (SOTA) methods on downstream applications with limited data resources. For instance, pre-trained models trained on large-scale, open-domain datasets have demonstrated SOTA performance in machine anomaly detection [21, 22]. In healthcare, pre-training techniques have been applied to detect conditions such as Alzheimer’s disease [23, 24] and heart failure [25]. Specifically for voice disorder analysis, speech pre-trained models have emerged as popular pre-trained audio feature extractors. Tirronen et al. explore the effectiveness of using wav2vec [26] feature extractor to detect voice disorders [27, 28]. Javanmardi et al. specifically targets the detection of the dysarthria speech using wav2vec models [29, 30]. Moreover, models pre-trained on image domain is also helpful for audio classification of voice pathologies [31]. Despite these advancements, audio-only approaches fail to incorporate vital visual information, which provides richer and more interpretable insights into vocal fold conditions, rendering them incapable of distinguishing left-sided from right-sided paralysis. Furthermore, the scarcity of clinical audio data poses a significant challenge for developing robust and generalizable machine learning models.
Visual modalities provide intuitive and interpretable features, enabling more effective clinical diagnoses. Foundational research established a strong correlation between the maximum separation of vocal folds and VFP [32], motivating the development of methods to quantitatively analyze this separation. For instance, Gloger et al. concentrated on segmenting the glottic region by extracting physical attributes such as shape and color [33]. Adamian et al. and Wang et al. utilized supervised learning techniques to track vocal fold movements and calculate the Anterior Glottic Angle (AGA), which measures the angular relationship between the two vocal folds [34, 35]. These studies introduced the Glottal Area Waveform (GAW) and the AGA waveform as diagnostic tools by computing the glottal area and AGA for each video frame. Pennington et al. expanded this line of research by incorporating additional features; however, their analysis predominantly relied on metrics derived from the AGA [36]. Despite these advances, current approaches face two key limitations: they primarily measure the separation between the left and right vocal folds, rendering them inadequate for distinguishing left-sided from right-sided paralysis, and they are restricted to pre-processed or brief video segments, often omitting sequences where vocal folds are not visible in raw endoscopic videos. These constraints hinder their application to raw, extended laryngoscopic videos commonly encountered in clinical practice, while also requiring manual annotation, which increases the workload for clinicians.
In this article, we introduce the Multimodal Laryngoscopic Video Analyzing System (MLVAS), a novel framework that leverages both audio and visual modalities to automatically process and analyze raw laryngeal videostroboscopic recordings. The proposed system comprises a multimodal front-end module, a feature extraction module, and a multimodal back-end classification module for detecting VFP. The front-end module identifies and extracts video segments that captures complete phonation cycles with clear vocal folds’ movement from raw endoscopic laryngeal videos. Subsequently, MLVAS extracts both audio and video features from these segments. By leveraging these complementary modalities, the system employs a classifier in the back-end for VFP detection. The primary contributions of this work are as follows:
@@
    We propose a multimodal system, MLVAS, that incorporates an audio-based Keyword Spotting (KWS) mechanism to automatically identify and extract relevant video segments containing complete phonation cycles and strobing sequences from raw laryngeal videostroboscopic recordings. This approach reduces the time required for clinicians to review lengthy video data.
@@
    To address the challenge posed by the scarcity of clinical data, the system leverages a pre-trained audio encoder for detecting vocal fold pathologies. To the best of our knowledge, this is the first application of a pre-trained audio model for a vocal fold paralysis prediction.
@@
    The proposed system further enhances VFP detection by incorporating visual features extracted from laryngeal video recordings with the audio features. Experimental results demonstrate that combining audio and video modalities leads to substantial improvements in VFP detection performance.
  • 4.
    The system introduces Left and Right Vocal Fold Dynamics (LVFDyn and RVFDyn) to reflect the vocal fold’s activity, enabling precise diagnosis of left- or right-sided VFP. Experimental findings highlight its effectiveness in diagnosing Unilateral Vocal Fold Paralysis (UVFP).
The organization of this paper is as follows: Section 2 introduces the proposed method for extracting key video segments, including audio keyword spotting, visual glottis detection, and strobing segment identification. Section 3 describes the methodology for audio feature extraction using a pre-trained audio encoder. Section 4 outlines the two-stage glottis image segmentation technique for video feature extraction, consisting of an initial U-Net-based segmentation followed by diffusion-based refinement. Additionally, this section introduces the computation of LVFDyn and RVFDyn features derived from the segmented glottis masks. Section 5 details the multimodal model for detecting both VFP and UVFP. Finally, Section 6 and Section 7 presents the experimental settings and results, and Section 8 concludes the paper.
Figure 1: The overview of our proposed MLVAS framework.
2 Multimodal Extraction of Key Video Segments
2.1 System Design
Our system aims to facilitate efficient clinical examinations by extracting key segments from laryngoscopic videos and providing objective indicators for specific laryngeal diseases. As Figure 1 shows, comprising two main modules – the voice module, and the video module – our system ensures the accurate observation of vocalization cycles and the clear visualization of the glottal area.
The voice module initially processes the audio extracted from the video using STFT to obtain spectrograms. Through the KWS technique, each frame is analyzed to detect patient vocalizations. This enables the preliminary segmentation of vocalization segments within the video.
Subsequently, the video module further refines these vocalization segments to obtain key video segments to form laryngoscopic highlights, ensuring the visibility of the vocal folds. Specifically, by utilizing the glottis object detection model, MLVAS can identify regions containing both the vocal folds and the glottis area in each frame. Moreover, given the importance of the stroboscopic portions in laryngoscopic videos for subjective analysis by physicians, we also include a stroboscopic video extraction method into our system.
@@ 2.1 System Design
2.2 Audio Processing Module
To get a better examination of the status of the vocal folds, patients are often pronouncing ”e” (although asked to pronounce ”a”) while the clinicians examine the larynx using the endoscope. In this way, the vibration of the vocal folds can be easily detected by the clinicians. Hence, by capturing the audio segments that pronounce ”e”, phonation cycles can be detected. To accomplish this, we developed a KWS model that is used to detect audio keywords in a sentence, such as ”Hey Siri”, ”Ok Google”, and in our case, the word ”e”.
The overall pipeline is illustrated in Figure 2. Initially, the input audio is transformed into a STFT spectrogram, converting the time-domain signal into a time-frequency representation for more effective analysis. The spectrogram is then segmented into chunks along the time axis, with each chunk containing a fixed number of frames. During the training phase, these chunks are randomly selected from each audio clip to ensure a diverse set of training samples. This method enhances the model’s generalization capabilities by incorporating spectrograms from various audio clips into a robust batch of training data.
@@ 2.2 Audio Processing Module
During inference, the input spectrogram is sliced into chunks using a sliding window mechanism. Each chunk is processed by the trained KWS model to generate decision results, as depicted in Figure 2. This KWS model can reliably detects the ”e” vocalization, providing critical prior knowledge for the subsequent analysis.
@@ 2.3 Video Processing Module
Although the audio KWS model is employed to detect the vocalization video segment of the phonation ”e”, these detected video segments may not always successfully capture the vocal folds and glottis on the visual side. To address this issue, we train a vocal fold detection model, utilizing the famous object detection model structure YOLO-v5 [38] on the BAGLS dataset [39], to further refine the time masks.
As depicted in Figure 1, our pipeline integrates a strobing video extraction module alongside the vocal fold image object detection module. This module is utilized to select the strobing segments of laryngeal videos by analyzing hue, saturation, and value (HSV) of video frames. The HSV parameters represent type, intensity, and brightness of colors, respectively. Transitions marked by empty frames typically occur at the beginning and end of strobing segments. We employ a unit step function to identify and mark all empty frames with a zero value in color (illustrated by the yellow line in Figure 3). This technique segments the video into several continuous non-empty frame sequences. Within these non-empty segments, we examine the variations in HSV values. The variations is calculated as the total number of changes within a time range. Due to the rapid color-changing characteristic of strobing videos (as shown in Figure 3), we are able to identify the strobing segments by calculating the frequency of HSV fluctuations within each non-empty segment, selecting the segment with the highest number of fluctuations. Suppose $v_{t-1},v_{t},v_{t+1}$ stands for the value at time step $t-1,t$ and $t+1$ respectively. Then, the frequency of HSV fluctuation within a time range of $T=[t_{0},t_{n}]$ is calculated as follows.
@@
3 Audio Modeling with Pretrained Models
The core idea of our method is to utilize large-scale pre-trained models as initialization and fine-tune them on clinical data for classification tasks. In our proposed system, we employ Dasheng [40], a state-of-the-art audio encoder, as the pre-trained model. Dasheng is trained on four extensive datasets encompassing speech, audio events, and music in a self-supervised manner. We selected Dasheng due to its large-scale dataset coverage, making it the most comprehensive open-source audio pre-trained model, to the best of our knowledge. Dasheng builds upon the principles of the Masked Audio Encoder (MAE) [41], which utilizes a transformer-based asymmetric encoder-decoder architecture. Unlike the traditional audio MAE [42], which masks and reconstructs time-frequency patches on the Mel-spectrogram, Dasheng applies masking along the time axis in a chunk-wise manner. The model incorporates learnable absolute positional embeddings, generates frame-level embeddings at a higher frequency, and processes consecutive chunks of Mel-spectrogram frames, enabling robust and efficient feature extraction for various kinds of downstream tasks.
During fine-tuning, we initialize the model with the pre-trained weights of Dasheng. As the encoded output of Dasheng is frame-wise, we compute the mean across all frame embeddings to derive an utterance-level embedding. A linear layer is attached in the end for final classification. It is noteworthy that we employ full fine-tuning in our experiments, optimizing all layers of Dasheng on the clinical dataset.
4 Visual Feature Extraction with Enhanced Glottis Image Segmentation